OpenStudy (anonymous):

y = mx + 6 how do you solve for m?

OpenStudy (anonymous):

to isolate m: y = mx + 6 -6 on both sides y-6 = mx /x on both sides (y-6)/x = m m = (y-6)/x <-- Final answer

OpenStudy (anonymous):

Say thank you and I can give you a medal, just respond

OpenStudy (anonymous):

thanks :)